-(b - 1) = 1 - b
True or false statement? If so, is there such a rule and I simply cannot find it? Thank you!

Answer by tommyt3rd(5050)

True. First notice that the negative sign means multiply the terms in the bracket by -1. Then the terms were rearranged and simplified:

-(b - 1) = (-1)b-(-1)(1)=-b-(-1)=-b+(-(-1))=-b+1=1+(-b)=1-b